# Explainable Malware Analysis: Concepts, Approaches and Challenges

### v1

@@ VI Conclusion
ML plays a crucial role in cybersecurity, yet these data-driven frameworks are susceptible to exploitation, misdirection, and circumvention, potentially leading to severe consequences. Explainability is essential for building trust in order to deploy models effectively in cybersecurity environments. Explainability techniques enhance the transparency of ML models and help identify critical features for malware detection, thus improving the capabilities of security analysts.
The field of XAI is steadily expanding. This paper discusses the taxonomy of XAI in malware analysis and reviews state-of-the-art approaches. We provide an in-depth examination of explainable malware classification and detection methods, summarizing the work of researchers to date. Our study systematically organizes various explainable malware-based approaches, making this information more accessible to researchers and others interested in this field.
We conclude the survey by identifying open research challenges and future directions in explainable malware analysis. Our findings highlight that significant work remains to be done in this area. This survey serves as a comprehensive guide for researchers exploring explainable malware detection, offering insights into the current landscape and stimulating research in unexplored domains within this dynamic and evolving field.

### v2

@@ VI Conclusion
ML-based techniques play a crucial role in cybersecurity, yet these data-driven frameworks are susceptible to exploitation, misdirection, and circumvention. Explainability is essential to enhance the transparency of these models and build trust in order to deploy them effectively for malware analysis.
This paper discusses explainable ML in malware analysis and reviews state-of-the-art approaches. We provide an in-depth examination of explainable malware classification and detection methods, summarizing the work of researchers to date. Our study systematically organizes various explainable malware-based approaches, making this information more accessible to researchers and others interested in this field.
We conclude the survey by identifying open research challenges and future directions in explainable malware analysis. This survey serves as a comprehensive guide for researchers exploring explainable malware detection, offering insights into the current landscape and stimulating research in unexplored domains within this dynamic and evolving field.